OpenStudy (anonymous):

The line through (-6,10) and (8,-6) also goes through the point (t,-1) for t=

hartnn (hartnn):

can you find the equation of line through those 2 given points ?

OpenStudy (anonymous):

well i know how to find the slope for (-6,10) and (8,-6) but i dont get whats with the other points

hartnn (hartnn):

ok, fin slope first

hartnn (hartnn):

then equation is just y-y1 = m (x-x1) x1, y1 is any one of those 2 points

OpenStudy (anonymous):

it would be y=-8/7x+22/7

hartnn (hartnn):

t,-1 will satisfy the equation of this line since it lies on it so, put x=t, y= -1 only t is unknown

OpenStudy (anonymous):

so it would end up as Y=-8/7x-8/7t-1?

hartnn (hartnn):

no...plug in y=-1 x=t y=-8/7x+22/7 -1 = -8/7 t +22/7

OpenStudy (anonymous):

ty

OpenStudy (tkhunny):

Alternatively, just looking for t and not finding the equation of the line. (-6,10) (8,-6) (t,-1) (10)(a) + (-6)(1-a) = 1 10a - 6 + 6a = 1 16a = 7 a = 7/16 Then: (-6)(7/16) + (8)(1 - 7/16) = t